what is 9 and 3/4 - 4 and 1/2 =
1
Expert's answer
2012-11-02T09:33:08-0400
9+(3/4 - 4)+1/2 =9+3/4 - 4+1/2 =5+(3+2)/4 =5+5/4=25/4